Margot Window Coverings Ltd. is preparing its annual profit plan. As a part of this process, budgeted material handling costs of $70,000 will be allocated to each of its product lines based on the following:

Draperies Blinds
Units produced 500 500
Material moves per product line 3 9
Direct labour hours per unit 140 140

Question 1: under a costing system that allocated overhead on the basis of direct labour-hours, what would the material handling costs assigned to 1 unit of draperies be?

a. $35

b. $70

c. $90

d. $100

Question 2: Under activity-based costing, where the cost driver is the material moves, what would be the material handling costs traced to one set of draperies?

a. $17.50

b. $35

c. $46.67

d. $70
Answer #1

Solution 1:

Material handling cost assigned to one unit of draperies = $70,000*140 / (1000*140) = $70 per unit

Hence option b is correct.

Solution 2:

Material handling cost assigned to draperies = $70,000 * 3/12 = $17,500

Material handling cost per unit of draperies = $17,500 / 500 = $35 per unit

Hence option b is correct.